OpenStudy (anonymous):

Can somebody help me please. 2(L+29)>=134

OpenStudy (anonymous):

\[2(L+29)\ge134\] \[2L+58 \ge 134\] \[2L \ge 76\] \[L \ge 38\]

OpenStudy (anonymous):

so this is how to find the possible value for the lenght

OpenStudy (anonymous):

This says that the length has to be greater than 38 or equal to 38

OpenStudy (anonymous):

ok because in my homework it tells me to find all possible values for the length of the deck

OpenStudy (anonymous):

is that the answer or the possible value for the lenght

OpenStudy (anonymous):

Anything 38 or greater is the possible length of the deck.

OpenStudy (anonymous):

thank you so much ,its easier when i put it in to words .thank you so much again

OpenStudy (anonymous):

No problem.